# Split and Rephrase with Large Language Models

### v2

@@ Abstract
The Split and Rephrase task, which consists in splitting complex sentences into a sequence of shorter grammatical sentences, while preserving the original meaning, can facilitate the processing of complex texts for humans and machines alike. In this work, we describe an approach based on large language models, which improves over the state of the art by large margins on all the major metrics for the task, on publicly available datasets. We also describe results from two human evaluations that further establish the significant improvements obtained with large language models and the viability of the approach. We evaluate different strategies, including fine-tuning pretrained language models of varying parameter size, and applying both zero-shot and few-shot in-context learning on instruction-tuned language models. Although the latter were markedly outperformed by fine-tuned models, they still achieved promising results overall. Our results thus demonstrate the strong potential of different variants of large language models for the Split and Rephrase task, using relatively small amounts of training samples and model parameters overall.
@@
[…]
@@ 1 Introduction
Transforming complex sentences into a sequence of shorter sentences, while preserving the original meaning, is an important task which can facilitate the processing of complex texts for humans and machines alike. It is, for instance, an important operation in text simplification Zhu et al. (2010); Narayan and Gardent (2014); Shardlow (2014), where simpler split sentences can provide content that is easier to read and comprehend.
@@
This task, commonly referred to as Split and Rephrase (SPRP) (Narayan et al., 2017), is typically performed via dedicated transformation models, e.g. sequence-to-sequence modelling with a copy mechanism (Aharoni and Goldberg, 2018) or graph-based neural segmentation (Gao et al., 2021). The main limitation of this type of approach is the dependency on training datasets of paired complex and simple sentences, which are scarce across languages and domains. As an alternative, Kim et al. (2021) exploit one-to-many alignments in parallel corpora and pivot machine translation (MT) to create an aligned corpus of complex and split sentences, which can be exploited with sequence-to-sequence models. For this type of approach, the main bottleneck is the dependency on parallel corpora with sufficient many-to-one alignments, which can also be scarce overall.
Pretrained large language models (LLMs), such as the GPT models (Radford et al., 2018) based on the Transformer architecture (Vaswani et al., 2017), have demonstrated their impressive potential to generate fluent text from natural language prompts, opening new research paths for a large number of downstream tasks. Their potential has been studied, for instance, for MT Bawden and Yvon (2023); Zhang et al. (2023), for translation quality estimation Kocmi and Federmann (2023), or for the generation of synthetic sentence similarity datasets via hand-crafted prompts Schick and Schütze (2021). We follow the spirit of these approaches and evaluate the potential of LLMs for the controlled generation of sequences of meaning-preserving short sentences from complex ones.
We measure the ability of different pretrained LLMs to perform the SPRP task, using task-specific prompts that take a complex sentence as input and probe the language model for a continuation in the form of split sentences. We evaluate our LLM-based approach on publicly available English datasets and explore controlled generation under zero shot, few shot in-context learning, and fine-tuning scenarios. Our analysis includes the use of LLMs with varying number of parameters, to establish minimal and optimal configurations for the SPRP task. We also evaluate the accuracy of LLMs trained only on the next token generation task and models further tuned for instruction completion. Finally, we measure the impact of training data size on SPRP accuracy of fine-tuned LLMs.
Overall, our results demonstrate the strong potential of pretrained LLMs for SPRP, outperforming the state of the art by large margins across multiple metrics. We also report results of two manual evaluations, a comparative 3-way ranking task and a qualitative evaluation of LLM-generated SPRP output, both of which further establish the gains and output quality achievable by our approach.
@@ 2 Related Work
The Split and Rephrase task was first proposed by Narayan et al. (2017), who created the WebSplit dataset from RDF subject-predicate-object semantic triples and also evaluated initial benchmarks using different types of models, including statistical MT models and encoder-decoder LSTMs with attention Bahdanau et al. (2015). Aharoni and Goldberg (2018) reported that the initial baselines suffered from data memorisation and lack of RDF relation types in the training set. They proposed a new data split and a new benchmark using LSTMs using attention and a copy mechanism (Gu et al., 2016), to bias the model towards copying the source. Botha et al. (2018) argued that WebSplit was limited by its small vocabulary and unnatural linguistic expressions, and presented WikiSplit, a corpus created using Wikipedia data that takes advantages of edits made on a page to split or fuse sentences. They showed that models trained with WikiSplit dataset obtained better BLEU Papineni et al. (2002) scores than those trained on WebSplit. However, they also noted that WikiSplit is a relatively noisy dataset, not suited for evaluation.
@@
In Das et al. (2018), simple sentences are generated from complex and compound sentences by leveraging dependency parsing information, achieving promising results. Niklaus et al. (2019) employed handcrafted rules, recursively applied to complex sentences to generate simpler ones, which they reported to be beneficial in downstream tasks. Recent approaches have exploited neural modelling, and in particular the Transformer architecture (Vaswani et al., 2017). Thus, Malmi et al. (2019) proposed a tagging approach using a BERT model (Devlin et al., 2019) as encoder and a single Transformer decoder layer. Stahlberg and Kumar (2020) also propose a tagging model, at the span level instead of the token level, obtaining improvements with a Pointer Network (Vinyals et al., 2015) for end of span prediction and a two-layer Transformer decoder. Berlanga Neto and Ruiz (2021) introduced a cross-lingual model based on part-of-speech tags, GRU sequence-to-sequence modelling, and BERT for next word prediction. Recently, AlAjlouni and Li (2023) reported improvements via transfer knowledge of data generated by the DISSIM rule-based system (Niklaus et al., 2019) to fine-tune an encoder-decoder T5 model .
@@
In this work, we compare our LLM-based approach with two recent approaches and associated corpora. First, we address the work of Kim et al. (2021), who proposed BiSECT, a method and corpus based on extracting 1-2 alignments from parallel bilingual corpora and machine translating the sentences into the same language. They showed that BiSECT was of higher quality than WikiSplit, and that sequence-to-sequence models trained on it achieved better results on the SPRP task overall. Additionally, we compare our work to Gao et al. (2021), who proposed the ABCD framework where SPRP is performed via neural graph segmentation, according to the number of predicates in the complex sentence, showing that this approach approached or outperformed the performance of parsing-based and encoder-decoder models. They also provided two new datasets for the SPRP task: MinWiki, a reduced version of WikiSplit where only pairs of complex and simple sentences with a matching number of predicates are kept, and DeSSE, a corpus of students’ opinion essays in social science consisting of complex sentences and corresponding splits; we used the latter in our experiments, in addition to BiSECT.
@@ 3 Methodology
Our approach is based on a simple idea: prompt a LLM to generate sequences of sentences from a single complex sentence. We explored two variants to control text generation for these tasks.
We first fine-tuned pretrained LLMs on task-specific prompts, querying for complex sentence split. The objective in this scenario is twofold: (i) measure the performance of LLMs that have not been previously fine-tuned for instruction-based task completion, and (ii) evaluate the accuracy of LLMs depending on their number of parameters. The latter point is relevant in particular to evaluate minimal and sufficient needs in terms of computational resources for the task.
Additionally, we evaluate scenarios where we directly prompt LLMs that have been previously tuned on instruction datasets. The goal in this case is to measure the ability of this type of readily trained models to perform the two tasks at hand, with or without positive examples. Under this scenario, we explore both zero-shot and in-context few-shot task completion, by directly prompting the model for completion in the first case, and providing a limited number of examples within the prompt in the latter case.
@@ 4.1 Corpora
To train and evaluate our models we selected two recent datasets in English for the SPRP task: DeSSE (Gao et al., 2021) and BiSECT (Kim et al., 2021). [^1] We use the datasets made available in the following repositories, as of April 2023: https://github.com/serenayj/DeSSE and https://github.com/mounicam/BiSECT. Statistics for each corpus are provided in Table 1, in terms of number of aligned pairs of complex and split sentences. We indicate the figures for the original dataset (Original) and the versions we employed in this work (Selected).
@@
For testing, we filtered all pairs of complex and split sentences in each corpus where the latter consisted of only one sentence, instead of multiples ones. There were 286 such instances in the DeSSE corpus, and 3 in BiSECT. We used the filtered version of the test datasets for all experiments reported below, as we aimed to evaluate model performance strictly on pairs of complex sentences and corresponding sequences of multiple sentences.
In terms of training data, we also filtered the DeSSE training set along the same lines, and used the resulting dataset for fine-tuning. For BiSECT, we randomly sampled the same amount of training data, for comparison purposes. As development data, since the original sets for DeSSE were rather small (42 examples), we extracted 2,000 pairs from the training set for both datasets.
The two corpora differ in several respects. First, in terms of size, BiSECT is orders of magnitude larger in terms of training and development datasets, although the filtered version of both test sets are of comparable size. In terms of domain, DeSSE contains students opinion essays collected from a social science class and is therefore less varied in terms of topics than BiSECT, which was created from a large variety of parallel corpora. Finally, the average length of the complex sentences, computed in terms of tokens over tokenised sentences, is also markedly different: for the training datasets, it amounts to an average of 40.14 tokens in BiSECT, contrasting with 25.25 in DeSSE. These differences are maintained for the test sets, with average lengths of 39.55 for BiSECT and 24.10 for DeSSE.
@@ 4.2 Models
As pretrained LLMs, we used the Pythia suite (Biderman et al., 2023), a collection of Transformer-based decoder-only autoregressive generative language models available in different parameter size and trained on the Pile dataset (Gao et al., 2020). We selected the following variants for our experiments, in terms of number of model parameters: 410M, 1B, 1.4B, 2.8B, 6.9B and 12B. We fine-tuned each of these models with LoRA (Hu et al., 2022), using default parameters and the following prompt, where $\alpha$ is a complex sentence and $\beta$ its corresponding split sentences, sampled from the selected portions of the DeSSE and BiSECT training sets: [^2] See Appendix A for more details on LLM training.
  • INSTRUCTION: Split the sentence below into separate sentences.
  • INPUT: $\alpha$
  • RESPONSE: $\beta$
Although other types of prompts are conceivable for the task, with different instructions or markers to denote the components of a prompt, preliminary results with prompt variants did not lead to significantly different results and we left further exploration on this matter for future research. At inference time, the response is empty and to be completed by the model.
@@
As indicated in Section 3, we used two different approaches to probe these model variants. For zero-shot probing, we used the prompt patterns described above, leaving the response empty for the model to complete. For in-context learning, we used a few-shot approach based on $n$ examples prepended to the prompt. Thus, each query is composed of $n$ concatenated prompts that follow the previously described pattern, along with the last prompt where the response is left empty. For our experiments, we selected the first 5 examples from the development set as examples, although other variants are of course possible.
Figure 1: Split and Rephrase results on the DeSSE (left) and BiSECT (right) test sets, with Pythia model variants fine-tuned over data from the DeSSE (-DE) or BiSECT (-BI) training data.
@@ 4.3 Evaluation
Following Kim et al. (2021), we used three separate metrics to evaluate our models. SARI (Xu et al., 2016), the standard simplification metric, and more specifically the version computed with the EASSE toolkit (Alva-Manchego et al., 2019) [^6] https://github.com/feralvam/easse, which provides the proportion of ngrams that are added (add in what follows), kept (keep) or deleted (del). As a second metric, we used the normalised BertScore (hereafter, BESC) (Zhang et al., 2020b), [^7] https://github.com/Tiiiger/bert_score which has proved successful to measure meaning preservation. BLEU scores, computed with sacreBLEU (Post, 2018), are reported as well, as a precision measure of n-gram overlap between hypotheses and references.
We also included an additional metric for the task to indicate output compliance (CPL), i.e. the percentage of output sequences that include at least two separate sentences, to directly identify unsplit output. This metric was computed with a simple approximation, counting the number of end of sentence punctuation marks in the output sentence, including period, question and exclamation.
We compare our results with those obtained using the BiSECT and ABCD approaches of Kim et al. (2021) and Gao et al. (2021), respectively, which represent the state of the art for the selected corpora. For the BiSECT approach, we used the publicly shared output available for their model. [^8] https://github.com/mounicam/BiSECT/tree/main/outputs/
bisect/our_model_bisect.txt
For the ABCD graph segmentation approach, we recomputed the output of their models using the publicly available code. [^9] https://github.com/serenayj/DeSSE

### v4

@@ Abstract
The Split and Rephrase (SPRP) task, which consists in splitting complex sentences into a sequence of shorter grammatical sentences, while preserving the original meaning, can facilitate the processing of complex texts for humans and machines alike. It is also a valuable testbed to evaluate natural language processing models, as it requires modelling complex grammatical aspects. In this work, we evaluate large language models on the task, showing that they can provide large improvements over the state of the art on the main metrics, although still lagging in terms of splitting compliance. Results from two human evaluations further support the conclusions drawn from automated metric results. We provide a comprehensive study that includes prompting variants, domain shift, fine-tuned pretrained language models of varying parameter size and training data volumes, contrasted with both zero-shot and few-shot approaches on instruction-tuned language models. Although the latter were markedly outperformed by fine-tuned models, they may constitute a reasonable off-the-shelf alternative. Our results provide a fine-grained analysis of the potential and limitations of large language models for SPRP, with significant improvements achievable using relatively small amounts of training data and model parameters overall, and remaining limitations for all models on the task.
@@
David Ponce∗1,2and Thierry Etchegoyhen∗1and Jesús Calleja1,2and Harritxu Gete1,2 1 Fundación Vicomtech, Basque Research and Technology Alliance (BRTA) 2 University of the Basque Country UPV/EHU {adponce,tetchegoyhen,jcalleja,hgete}@vicomtech.org
@@ 1 Introduction
Transforming complex sentences into a sequence of shorter sentences, while preserving the original meaning, can facilitate the processing of complex texts for humans and machines alike. It is, for instance, an important operation in text simplification Zhu et al. (2010); Narayan and Gardent (2014); Shardlow (2014), where simpler split sentences can provide content that is easier to read and comprehend. This task, commonly referred to as Split and Rephrase (SPRP) (Narayan et al., 2017), also provides a relevant framework to evaluate natural language processing models, as complex linguistic properties need to be tackled to provide optimal splits and appropriate rephrasing.
@@
The SPRP task is typically performed via dedicated transformation models, e.g. sequence-to-sequence modelling with a copy mechanism (Aharoni and Goldberg, 2018) or graph-based neural segmentation (Gao et al., 2021). The main limitation for this type of approach is the dependency on training datasets of paired complex and split sentences, which are scarce across languages and domains. As an alternative, Kim et al. (2021) exploit one-to-many alignments in parallel corpora and pivot machine translation (MT) to create an aligned corpus of complex and split sentences, which can in turn be exploited by sequence-to-sequence models. For this type of approach, the main bottleneck is the dependency on parallel corpora with sufficient many-to-one alignments and on quality pivot machine translation models.
Large language models (LLMs), such as the GPT models (Radford et al., 2018) based on the Transformer architecture (Vaswani et al., 2017), have demonstrated their strong potential on a large number of downstream tasks (Radford et al., 2019; Brown et al., 2020). In this work, we measure their performance on the SPRP task, prompting the models to generate split and rephrase hypotheses from complex sentences. We evaluate different variants of the approach on datasets of varying complexity, contrasting zero shot, few shot in-context learning, and fine-tuning scenarios. We notably measure the impact of parameter and training data size, domain shift, and prompt optimisation.
Overall, our results demonstrate that LLMs outperform the previous state of the art by large margins across multiple metrics, with remaining relative deficiencies, notably in terms of split generation. Results from two human evaluations, a comparative 3-way ranking task and a qualitative evaluation, further support our conclusions. Our main contributions can be summarised as follows:
  • New state of the art on the SPRP task, established on publicly available datasets of varying complexity, over multiple metrics and two human evaluations.
  • A comprehensive evaluation of LLMs for SPRP, covering LLM variants with and without instruction tuning, along with the impact of different prompting strategies, parameter size, training data volumes, and domain shift.
  • Empirical results on LLM strengths and limitations for SPRP, in particular over splitting compliance and rephrasing variability.
@@ 2 Related Work
The Split and Rephrase task was first proposed by Narayan et al. (2017), who created the WebSplit dataset from RDF semantic tuples and provided benchmarks with statistical MT models and LSTMs with attention. Aharoni and Goldberg (2018) identified data representation in this dataset and provided enhancements along with a new benchmark augmenting the LSTM with a copy mechanism (Gu et al., 2016). To address some limitations of WebSplit, notably in terms of vocabulary and naturality, Botha et al. (2018) presented WikiSplit, a corpus created using Wikipedia data that takes advantages of edits made on a page to split or fuse sentences. They showed that models trained on WikiSplit improved those trained on WebSplit in terms of BLEU Papineni et al. (2002). However, they also noted that WikiSplit is a relatively noisy dataset, not suited for evaluation.
@@
Among other approaches, Das et al. (2018) leveraged dependency parsing information for the SPRP task, whereas Niklaus et al. (2019) employed handcrafted rules, recursively applied to complex sentences to generate simpler ones. Recent approaches have exploited neural modelling, in particular the Transformer architecture. Thus, Malmi et al. (2019) proposed a tagging approach using a BERT model (Devlin et al., 2019) as encoder and a single Transformer decoder layer. Stahlberg and Kumar (2020) also explored a tagging model, at the span level instead of the token level, achieving improvements with Pointer Networks (Vinyals et al., 2015) for end of span prediction and a two-layer Transformer decoder. Berlanga Neto and Ruiz (2021) introduced a cross-lingual model based on part-of-speech tags, GRU sequence-to-sequence modelling, and BERT for next word prediction. Recently, AlAjlouni and Li (2023) reported improvements via transfer knowledge of data generated by the DISSIM rule-based system (Niklaus et al., 2019) to fine-tune an encoder-decoder T5 model .
@@
In this work, we compare our LLM-based variants with two recent approaches and associated corpora. First, we address BiSECT (Kim et al., 2021), a method and corpus based on extracting 1-2 alignments from parallel bilingual corpora and pivot machine translation. BiSECT achieved higher quality than WikiSplit, and sequence-to-sequence models trained on the corpus improved on the SPRP task. Additionally, we compare our work to Gao et al. (2021), who proposed the ABCD framework where SPRP is performed via neural graph segmentation, according to the number of predicates in the complex sentence, approaching or outperforming parsing-based and encoder-decoder models. They also provided two new datasets for the SPRP task: MinWiki, a reduced version of WikiSplit where only pairs of complex and simple sentences with a matching number of predicates are kept, and DeSSE, a corpus of students’ opinion essays in social science consisting of complex sentences and corresponding splits; we used the latter in our experiments, in addition to BiSECT.
@@ 3 Methodology
The overall approach to SPRP with LLM is simple: probe a LLM to generate sequences of meaning-preserving sentences from a single complex sentence. We explored two main variants for the task.
We first fine-tuned pretrained LLMs on pairs of complex and split sentences. The objective in this scenario is twofold: (i) measure the performance of bare LLMs on the SPRP task, and (ii) evaluate LLM accuracy with varying number of parameters. For these models, we also evaluate different approaches to prompting, comparing soft and manually crafted prompts.
Additionally, we evaluate scenarios where we directly prompt LLMs that have been previously tuned on instruction datasets. The goal in this case is to measure the ability of this type of readily-trained models to perform the two tasks at hand, with or without positive examples. Under this scenario, we explore both zero-shot and in-context few-shot task completion, by directly prompting the model for completion in the first case, and providing a limited number of examples within the prompt in the latter case.
@@ 4.1 Corpora
To train and evaluate our models we first selected the aforementioned English datasets: DeSSE and BiSECT. [^1] We use the datasets made available in the following repositories, as of April 2023: https://github.com/serenayj/DeSSE and https://github.com/mounicam/BiSECT. Statistics for each corpus are provided in Table 1, in terms of number of pairs of complex and split sentences. We indicate the figures for the original dataset (Original) and the versions we employed in this work (Selected).
We excluded all pairs where the split sentence did not contain at least one inner end-of-sentence punctuation mark (286 instances in DeSSE and 3 in BiSECT). For training, we down-sampled BiSECT to the size of DeSSE for comparison purposes. As development data, since the original sets for DeSSE were rather small (42 examples), we sampled 2,000 pairs from both training datasets.
@@
We chose these two corpora as they present different challenges for SPRP. In terms of topics, DeSSE is based on social science opinion essays and is therefore less varied than BiSECT, which was created from a large variety of parallel corpora. In terms of complex sentence average length, the two datasets are also markedly different: for training, the average is 40.14 tokens in BiSECT, contrasting with 25.25 in DeSSE; in the test sets, the averages were 39.55 and 24.10, respectively.
Finally, for our qualitative manual evaluation, we used the Wiki-BM corpus Zhang et al. (2020a), as it was manually crafted for the SPRP task.
@@ 4.2 Models
As pretrained LLMs, we used the Pythia suite (Biderman et al., 2023), a collection of Transformer-based decoder-only autoregressive generative language models available in different parameter sizes and trained on the Pile dataset (Gao et al., 2020). We selected the following variants for our experiments, in terms of number of model parameters: 410M, 1B, 1.4B, 2.8B, 6.9B and 12B.
@@
Prompting.
We used different prompting strategies, depending on the model variant. For bare LLMs, we first compare the use of a simple hand-crafted prompt, namely:
  • INSTRUCTION: Split the sentence below into separate sentences.
  • INPUT: $\alpha$.
  • RESPONSE: $\beta$
, where $\alpha$ and $\beta$ are provided at training time and the model is tasked to complete $\beta$ at inference time. For this prompting approach, we fine-tuned the Pythia models with LoRA (Hu et al., 2022), using default parameters (see Appendix A) .
Additionally, we experimented with soft-prompting methods, which have been shown to improve LLM performance on several tasks, addressing the limitations of manual prompt design. We selected two such approaches, namely prompt-tuning Lester et al. (2021) and P-tuning Liu et al. (2023). In both cases, trainable prompt embeddings are added to a frozen base model, with P-tuning also involving a dedicated encoder.
For zero-shot probing, we used the hand-crafted prompt described above, leaving the response empty for the model to complete. For in-context learning, we used a few-shot approach based on $n$ examples prepended to the prompt. Thus, each query is composed of $n$ concatenated prompts that follow the previously described pattern, along with the last prompt where the response is left empty. For our experiments, we selected the first 5 examples from the development set as examples, although other variants are of course possible.
Preliminary experiments with alternative zero-shot prompt variants led to little variation in results across metrics, with our selected prompt achieving more balanced results across metrics. For few-shot sampling, larger sets of 10 examples provided minor improvements on some metrics, but degradation in terms of compliance. Results with alternative zero-shot and few-shot prompting are provided in Appendix C.
@@ 4.3 Evaluation
Following Kim et al. (2021), we used three separate metrics to evaluate our models. SARI (Xu et al., 2016), the standard simplification metric as computed by the EASSE toolkit (Alva-Manchego et al., 2019), [^5] https://github.com/feralvam/easse which provides the proportion of n-grams that are added (add in what follows), kept (keep) or deleted (del). As a second metric, we used the normalised BertScore (hereafter, BESC) (Zhang et al., 2020b), [^6] https://github.com/Tiiiger/bert_score which has proved successful to measure meaning preservation. BLEU scores, computed with sacreBLEU (Post, 2018), are reported as well, as a precision measure of n-gram overlap between hypotheses and references.
We also included an additional metric for the task to indicate output compliance (CPL), i.e. the percentage of output sequences that include at least two separate sentences, to directly identify unsplit output. This metric was computed with a simple approximation, counting the number of end of sentence punctuation marks in the output sentence period, question and exclamation marks.
We compare our results with those obtained using the BiSECT and ABCD approaches of Kim et al. (2021) and Gao et al. (2021), respectively, the state of the art for the selected corpora. For BiSECT, we used their publicly shared output directly, filtering the sentences that were excluded from the Selected test set described in Section 4.1, to maintain strict comparability with the other tested methods. [^7] https://github.com/mounicam/BiSECT/tree/main/outputs/ For ABCD, the output was computed with the publicly shared models and code, [^8] https://github.com/serenayj/DeSSE on the Selected test subset, for a fair comparison.
Figure 1: Impact of parameter size on the DeSSE (left) and BiSECT (right) test sets, with Pythia model variants fine-tuned over data from the DeSSE (-DE) or BiSECT (-BI) training data.
@@ 5 SPRP with Large Language Models
5.1 Prompting Variants
Table 2 shows the results of soft-prompting and LoRA fine-tuning. Excepting two BertScore results, where P-Tuning was marginally better, LoRA adaptation with our hand-crafted simple prompt achieved the best results overall. The differences were relatively small between the three approaches in terms of SARI and BESC, except for PRT-BI on BiSECT with significant losses on all metrics.
Model Test BESC SARI CPL
PRT-DE DeSSE 86.81 61.02 71.03
PTG-DE 87.67 63.29 72.82
LoRA-DE 88.78 69.47 94.64
PRT-DE BiSECT 73.42 27.03 63.10
PTG-DE 73.77 28.62 76.55
LoRA-DE 72.79 30.85 93.45
PRT-BI DeSSE 81.95 37.42 6.55
PTG-BI 82.00 40.20 15.67
LoRA-BI 80.54 48.80 71.43
PRT-BI BiSECT 37.00 1.55 8.79
PTG-BI 74.71 25.32 22.07
LoRA-BI 76.12 38.93 81.55
Table 2: Results for Prompt-tuning (PRT), P-Tuning (PTG) and LoRA adaptation with the Pythia-6.9B model trained on DeSSE (-DE) or BiSECT (-BI) data.
The most notable result are the significantly larger differences in terms of compliance, where LoRA was markedly better in all cases and both soft-prompting methods failed significantly overall, drastically so when trained on BiSECT data. The poor performance of both methods on the more complex SPRP data is noteworthy, considering the gains achieved with soft-prompting in other downstream tasks. Considering these results, we selected the more robust LoRA approach as our main adaptation method in the remaining experiments.